# Sun-Shine: A Large Language Model for Tibetan Culture

### v2

@@ Abstract
Tibetan, a minority language in China, features a highly intricate grammatical structure, characterized by four verb tenses and a tense system with frequent irregularities, contributing to its extensive inflectional diversity. Recently, advances in Large Language Models (LLMs) have transformed the paradigm in many domains. Despite the success in other fields, current LLMs often fall short in catering to the needs of domain experts like Tibetans, and the potential of LLMs for Tibetan culture is under-explored. The intrinsic reasons are the immense and intricate nature of Tibetan culture as well as the necessity for higher granularity and richness in knowledge. Simultaneously, the complexity and uniqueness of its grammatical structure, coupled with its status as a minority ethnic language, contribute to data scarcity, which remains a fundamental challenge. To alleviate these issues, we introduce Llama-Sunshine (Sun-Shine), the first large language model for Tibetan culture, which is expert in various Tibetan language processing tasks. Sun-Shine incorporates state-of-the-art model architectures optimized for Tibetan’s linguistic features. We also propose TIB-STC, a comprehensive dataset comprising diverse Tibetan texts such as literature, religious scripts, news, and conversational data, which is also the first large-scale dataset for Tibetan culture. Though comprehensive experiments, Sun-Shine not only demonstrates a higher level of knowledge expertise for Tibetan culture but also gains preliminary embodied intelligence capabilities in Tibetan language processing tasks, like language modeling, text classification, machine translation, and syntactic analysis. Moreover, it excels in low-resource scenarios, showcasing strong generalization capabilities. Code and data are avaliable: https://github.com/Vicentvankor/sun-shine.
@@ 1 Introduction
Large language models (LLMs) have gained significant attention for their remarkable performance across a wide range of natural language tasks, particularly following the release of ChatGPT Minaee et al. (2024), Gemini Team et al. (2023), Qwen-2.5 Yang et al. (2024a; b), Llama Touvron et al. (2023); Dubey et al. (2024) and DeepSeek Bi et al. (2024); Guo et al. (2025). Compared to smaller models, LLMs demonstrate superior generalization across a wide range of natural language processing (NLP) tasks and exhibit unique emergent abilities to solve previously unseen or complex challenges Achiam et al. (2023). Although ChatGPT remains non-open source, open-source communities have actively developed several alternative models with relatively affordable training costs, like LLaMA Touvron et al. (2023) and DeepSeek Bi et al. (2024); Guo et al. (2025), which establishes LLMs as promising solutions for real-world applications that demand effective communication and reasoning capabilities.
@@
Despite their many strengths, LLMs are not tailored to specific fields. Pre-training is the strategy that make LLMs become ”experts” in specific domains where LLMs need accurate and domain-specific expert knowledge, such as medical domain Wang et al. (2023a); Zhou et al. (2023). Also, there are some biases against LLMs GLM et al. (2024); Wang et al. (2023b); Blodgett et al. (2020). LLMs often reflect biases in training data, including gender, racial, and cultural biases, regional imbalances favoring English and Western cultures, and uneven domain coverage Achiam et al. (2023). For instance, GPT-3’s generated content was found to contain racial stereotypes in multiple experiments Brown et al. (2020). Mitigation requires diverse data, bias detection tools, and model adjustments to ensure fairness and reliability. Overall, data scarity is the key issue for these areas. Especially for some niche languages, such as Tibetan, data scarcity is the most direct and fundamental problem Liang et al. (2024). In addition, Tibetan itself has a highly complex structure, which is further compounded by the intricate syntax and grammar of the language Khysru et al. (2021). Assembling a dataset about such complex but elegant words is already a monumental task, but creating one that is fair and unbiased requires even greater effort. Only then can we develop a more specialized and unbiased LLM model on this basis.
@@
To address this issue, we introduce the TIB-STC, the first comprehensive Tibetan corpus dataset designed to address the issue of data scarcity in Tibetan language research. This dataset spans diverse domains, including law, public welfare, Tibetan literature, and Tibetan medicine, making it the first large-scale, detailed, and comprehensive Tibetan dataset in the world. On this basis, we proposed Llama-Sunshine (Sun-Shine), the large language model of Tibetan culture. the first LLM specifically tailored for Tibetan language processing. During its development, we incorporated Chinese translations and manual annotation techniques to improve training accuracy and enhance its performance.
@@
Through extensive experiments, Sun-Shine demonstrates superiority for Tibetan culture, even slightly better tham DeepSeek-R1 Guo et al. (2025) totally, showing a higher level of knowledge expertise in Tibetan Culture. Note that TIB-STC is based on criteria manually evaluated by Tibetan experts, and can accurately reflect the capabilities that LLMs possess in the field of Tibetan culture. Together, Sun-Shine and TIB-STC represent not only pioneering efforts as the first LLM and dataset for Tibetan but also significant advancements in addressing the broader challenges of building LLMs for minority languages and mitigating bias in underrepresented linguistic contexts.
@@
  • We introduce TIB-STC, the first large-scale Tibetan dataset, addressing data scarcity in Tibetan language research. Covering domains like law, welfare, literature, and medicine, it serves as a vital resource for advancing Tibetan NLP.
  • Experimental results demonstrate that Sun-Shine, trained and fine-tuned on our dataset, TIB-STC, significantly outperforms other LLMs in Tibetan language processing totally, demonstrating superiority for Tibetan culture.
@@ 2.1 Tibetan Language Processing
LLMs for rare or low-resource languages are relatively uncommon Gurgurov et al. (2024). The absence of corpus data is the primary and most critical issue Brown et al. (2020); Liang et al. (2024) like Tibetan. However, Tibetan-speaking population in Tibet represents a minority within China, with an even smaller subset capable of reading and writing the language proficiently Institute (2014). As a result, digital Tibetan language resources are exceptionally scarce. This scarcity extends to linguistic corpora, which are critical for the development and training of artificial intelligence models, thereby hindering advancements in AI, like LLMs, for Tibetan language processing Liang et al. (2024); Khysru et al. (2021).
For Tibetan itself, its syllables are two-dimensionally combined by Tibetan characters and we have sorted out the composition distribution of Tibetan characters, shown in Appendix Figure 7. As shown in Appendix Figure 6, the native Tibetan syllable structure, includes seven kinds of letters. There is a longitudinal structure in syllables which is called the upright unit, consisting of root letter, superscript letter, subscript letter, and vowel. Such complexity of Tibetan grammar and structure, combined with its status as a minority language, results in a severe scarcity of Tibetan data. To solve this issue, we propose the first large-scale Tibetan dataset, TIB-STC, the largest and the most comprehensive one in the world now, advancing Tibetan NLP and offering a foundation for future research and potential applications while serving as a model for developing NLP resources for other minority languages.
@@ 2.2 Large Language Model
LLMs, including GPT Achiam et al. (2023), Gemini Team et al. (2023), and ChatGLM GLM et al. (2024); Wang et al. (2023b), are highly parameterized models with billions of parameters, pre-trained on massive natural language datasets Chiang & Lee (2023). Their ability to excel in unseen tasks, given only task-specific instructions, underscores their versatility and advanced capabilities. By applying task-specific instructions and fine-tuning techniques Hu et al. ; Rafailov et al. (2024); Yu et al. (2021), LLMs can be optimized for specialized fields, enhancing their performance and domain expertise Zhou et al. (2023).
@@ 3 TIB-STC and Sun-Shine
3.1 Dataset: TIB-STC
TIB-STC, is the first comprehensive and detailed Tibetan corpus dataset for LLM, consisting of five categories of data as shown in the dataset branch of Table 1. Corpus is used for LLM pre-training, while other datasets focus on optimization and enhancement. For these four sub-datasets, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improves LLM’s alignment ability. Pre-Training involves the largest dataset in terms of instances, tokens, and storage, as it focuses on general language model training. In the stage of Supervised Fine-Tuning (SFT), datasets are smaller but still sizable, with a focus on task-specific fine-tuning (e.g., instructions and safety). Direct Preference Optimization (DPO) uses the smallest datasets but focuses on high-quality preference data for alignment and optimization.
Dataset Instance Tokens Storage Size
@@ 3.1 Dataset: TIB-STC
Table 1: The subcategories of TIB-STC and their data size statistics.
For TIB-STC, all data selection and cleaning are conducted by experts in the Tibetan language. Unlike modern Tibetan, this dataset preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics.
@@ 3.1.1 Corpus
Corpus for Sun-Shine’s pre-training consists of literature (66%), web pages (24%), and media (10%). Content distribution is led by philosophy (30%), medicine (24%), and culture (18%), with smaller portions from academic papers (10%), news (9%), and law (9%). Incorporating book texts enhances linguistic depth and strengthens long-term semantic modeling in LLMs. There are 122,756,184 such instances. Furthermore, leveraging large-scale web text data during pre-training facilitates the acquisition of diverse linguistic patterns and nuances, thereby improving the models’ proficiency in both natural language comprehension and generation. Finally, the inclusion of media data offers the unique advantage of establishing an expansive and reliable media knowledge base. As shown in Appendix Figure 8, we present a news-based training corpus where Chinese and Tibetan serve as the source languages, and English is provided as the translated language. Some of the original Chinese versions and Tibetan texts retain classical Chinese sentence structures, which are different from modern Chinese grammar.
@@ 3.1.2 Alpaca-Ti
Alpaca-Ti is used for instruction fine-tuning. We draw inspiration from Alpaca Taori et al. (2023a; b) project at Stanford University and develop the first Tibetan dataset specifically designed for self-supervised instruction fine-tuning in the world. It covers the following topics as shown in Table 2.
@@
Table 2: Details of Alpaca-Ti.
Alpaca-Ti encompasses a diverse array of topics and also has a longer length and better instruction fine-tuning performance, designed to enhance LLM’s generalization capabilities. It has a total of 42,676 instances. One data sample is shown in Appendix Figure 9.
@@ 3.1.3 Safety-Prompts-Ti
Safety-Prompts-Ti aims to improve the safety of large models on two major categories: Instruction Attacks and Typical Safety Scenarios, as shown in Table 3, and the data size is 38,156. LLM’s representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts Chao et al. (2024); Zheng et al. (2024); Sun et al. (2023). Human evaluation is a valuable approach; however, due to the extensive workload it entails, we opted to create a dataset to simulate the process in advance. One data sample is shown in Appendix Figure 10.
@@
Table 3: Statistical details of Instruction Attacks and Typical Safety Scenarios.
@@ 3.1.4 CValues-Ti
CValues-Ti is the dataset, used to improve the security of Tibetan language models. Considering work Moore et al. (2024); Rozen et al. (2024); Xu et al. (2023), we propose this branch of a small dataset because LLM can pose risks or has negative social impacts. For CValues-Ti, it has three categories (safe & reponsibility = rejection & positive advice; safe = rejection; unsafe = risk response). Then, under the same prompt, various types of responses can be categorized into positive and negative samples, differentiated by their levels of difficulty, so that there are three common combinations (safe + unsafe; safe & reponsibility + unsafe; safe& reponsibility + safe). Unlike the work Xu et al. (2023), We streamlined and condensed the instructions, and compared to their dataset (20K), our sub-dataset, CValues-Ti (2K), achieved equivalent results in subsequent experiments. One data sample is shown in Appendix Figure 11.
@@
hh-rlhf-Ti is also the first dataset that improves the model effect by introducing human intervention. It consists of two parts: ”harmless (23,822)” and ”helpful (23,589)” just like RLHF Bai et al. (2022); Kirk et al. , but we reset and summarize the previous dataset based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted prior to training for other objectives. This approach could be important in the future to prevent the development of intermediate, unaligned AI systems for Tibetan. It has a total of 47,411 instances, and one data sample is shown in Appendix Figure 12.
@@ 3.2 Sun-Shine
Sun-Shine is built on the LLaMA 3 Touvron et al. (2023); Dubey et al. (2024) architecture, a collection of multilingual base models with 7 to 65 billion parameters, openly available to the research community. For more accessible training, we adopt the LLaMA 3.1 (8B) model Dubey et al. (2024). Additionally, we integrate NLP model translation with human annotation Üstün et al. (2024) as supplementary data to enhance Sun-Shine’s accuracy.

### v3

# Sun-Shine: A Foundation Large Language Model for Tibetan Culture and Heritage

@@ Abstract
Tibetan, a minority language in China, features a highly intricate grammatical structure, characterized by four verb tenses and a tense system with frequent irregularities, contributing to its extensive inflectional diversity. Recently, advances in Large Language Models (LLMs) have transformed the paradigm in many domains. Despite the success in other fields, current LLMs often fall short in catering to the needs of domain experts like Tibetans, and the potential of LLMs for Tibetan culture is under-explored. The intrinsic reasons are the immense and intricate nature of Tibetan culture as well as the necessity for higher granularity and richness in knowledge. Simultaneously, the complexity and uniqueness of its grammatical structure, coupled with its status as a minority ethnic language, contribute to data scarcity, which remains a fundamental challenge. To alleviate these issues, we introduce Llama-Sunshine (Sun-Shine), the first large language model for Tibetan culture, which is expert in various Tibetan language processing tasks. Sun-Shine incorporates state-of-the-art model architectures optimized for Tibetan’s linguistic features. We also propose TIB-STC, a comprehensive benchmark comprising diverse Tibetan texts such as literature, religious scripts, news, and conversational data, which is also the first large-scale benchmark for Tibetan culture. Though comprehensive experiments, Sun-Shine not only demonstrates a higher level of knowledge expertise for Tibetan culture but also gains preliminary embodied intelligence capabilities in Tibetan language processing tasks, like language modeling, text classification, machine translation, and syntactic analysis. Moreover, it excels in low-resource scenarios, showcasing strong generalization capabilities.
Index Terms:
Large Language Model, Foundation Model, Tibetan Language Benchmark, Tibetan Language Processing
@@ 1 Introduction
Large language models (LLMs) have gained significant attention for their remarkable performance across a wide range of natural language tasks, particularly following the release of ChatGPT [1], Gemini [12], Qwen-2.5 [34, 33], Llama [2, 20] and DeepSeek [37, 38]. Compared to smaller models, LLMs demonstrate superior generalization across a wide range of natural language processing (NLP) tasks and exhibit unique emergent abilities to solve previously unseen or complex challenges [4]. Although ChatGPT remains non-open source, open-source communities have actively developed several alternative models with relatively affordable training costs, like LLaMA [2] and DeepSeek [37, 38], which establishes LLMs as promising solutions for real-world applications that demand effective communication and reasoning capabilities.
@@
Despite their many strengths, LLMs are not tailored to specific fields. Pre-training is the strategy that make LLMs become ”experts” in specific domains where LLMs need accurate and domain-specific expert knowledge, such as medical domain [3, 5]. Also, there are some biases against LLMs [6, 7, 8]. LLMs often reflect biases in training data, including gender, racial, and cultural biases, regional imbalances favoring English and Western cultures, and uneven domain coverage [4]. For instance, GPT-3’s generated content was found to contain racial stereotypes in multiple experiments [9]. Mitigation requires diverse data, bias detection tools, and model adjustments to ensure fairness and reliability. Overall, data scarity is the key issue for these areas. Especially for some niche languages, such as Tibetan, data scarcity is the most direct and fundamental problem [11]. In addition, Tibetan itself has a highly complex structure, which is further compounded by the intricate syntax and grammar of the language [10]. Assembling a benchmark about such complex but elegant words is already a monumental task, but creating one that is fair and unbiased requires even greater effort. Only then can we develop a more specialized and unbiased LLM model on this basis.
@@
To address this issue, we introduce the TIB-STC, the first comprehensive Tibetan corpus benchmark designed to address the issue of data scarcity in Tibetan language research. This benchmark spans diverse domains, including law, public welfare, Tibetan literature, and Tibetan medicine, making it the first large-scale, detailed, and comprehensive Tibetan benchmark in the world. On this basis, we proposed Llama-Sunshine (Sun-Shine), the large language model of Tibetan culture. the first LLM specifically tailored for Tibetan language processing. During its development, we incorporated Chinese translations and manual annotation techniques to improve training accuracy and enhance its performance.
@@
Through extensive experiments, Sun-Shine demonstrates superiority for Tibetan culture, even slightly better tham DeepSeek-R1 [38] totally, showing a higher level of knowledge expertise in Tibetan Culture. Note that TIB-STC is based on criteria manually evaluated by Tibetan experts, and can accurately reflect the capabilities that LLMs possess in the field of Tibetan culture. Together, Sun-Shine and TIB-STC represent not only pioneering efforts as the first LLM and benchmark for Tibetan but also significant advancements in addressing the broader challenges of building LLMs for minority languages and mitigating bias in underrepresented linguistic contexts.
@@
  • We introduce TIB-STC, the first large-scale Tibetan benchmark, addressing data scarcity in Tibetan language research. Covering domains like law, welfare, literature, and medicine, it serves as a vital resource for advancing Tibetan NLP.
  • Experimental results demonstrate that Sun-Shine, trained and fine-tuned on our benchmark, TIB-STC, significantly outperforms other LLMs in Tibetan language processing totally, demonstrating superiority for Tibetan culture.
@@ 2.1 Tibetan Language Processing
LLMs for rare or low-resource languages are relatively uncommon [13]. The absence of corpus data is the primary and most critical issue [9, 11] like Tibetan. However, Tibetan-speaking population in Tibet represents a minority within China, with an even smaller subset capable of reading and writing the language proficiently [16]. As a result, digital Tibetan language resources are exceptionally scarce. This scarcity extends to linguistic corpora, which are critical for the development and training of artificial intelligence models, thereby hindering advancements in AI, like LLMs, for Tibetan language processing [11, 10, 26, 19].
For Tibetan itself, its syllables are two-dimensionally combined by Tibetan characters and we have sorted out the composition distribution of Tibetan characters, shown in Fig. 1.
Figure 1: Characters Belong to Each Component
As shown in Fig. 2, the native Tibetan syllable structure, includes seven kinds of letters. There is a longitudinal structure in syllables which is called the upright unit, consisting of root letter, superscript letter, subscript letter, and vowel. Such complexity of Tibetan grammar and structure, combined with its status as a minority language, results in a severe scarcity of Tibetan data. To solve this issue, we propose the first large-scale Tibetan benchmark, TIB-STC, the largest and the most comprehensive one in the world now, advancing Tibetan NLP and offering a foundation for future research and potential applications while serving as a model for developing NLP resources for other minority languages.
Figure 2: One Example of the Native Tibetan Syllable Structure
@@ 2.2 Large Language Model
LLMs, including GPT [4], Gemini [12], and ChatGLM [6, 7], are highly parameterized models with billions of parameters, pre-trained on massive natural language benchmarks [14]. Their ability to excel in unseen tasks, given only task-specific instructions, underscores their versatility and advanced capabilities. By applying task-specific instructions and fine-tuning techniques [15, 17, 18], LLMs can be optimized for specialized fields, enhancing their performance and domain expertise [5].
@@ 3 TIB-STC and Sun-Shine
3.1 Benchmark: TIB-STC
TIB-STC, is the first comprehensive and detailed Tibetan corpus benchmark for LLM, consisting of five categories of data as shown in the benchmark branch of Table I.
TABLE I: Subcategories of TIB-STC and Data Size Statistics
Benchmark Instance Tokens Storage Size
@@ 3.1 Benchmark: TIB-STC
Corpus is used for LLM pre-training, while other benchmarks focus on optimization and enhancement. For these four sub-benchmarks, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improves LLM’s alignment ability. Pre-Training involves the largest benchmark in terms of instances, tokens, and storage, as it focuses on general language model training. In the stage of Supervised Fine-Tuning (SFT), benchmarks are smaller but still sizable, with a focus on task-specific fine-tuning (e.g., instructions and safety). Direct Preference Optimization (DPO) uses the smallest benchmarks but focuses on high-quality preference data for alignment and optimization.
For TIB-STC, all data selection and cleaning are conducted by experts in the Tibetan language. Unlike modern Tibetan, this benchmark preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics.
@@ 3.1.1 Corpus
Corpus for Sun-Shine’s pre-training consists of literature (66%), web pages (24%), and media (10%). Content distribution is led by philosophy (30%), medicine (24%), and culture (18%), with smaller portions from academic papers (10%), news (9%), and law (9%). Incorporating book texts enhances linguistic depth and strengthens long-term semantic modeling in LLMs. There are 122,756,184 such instances.
Furthermore, leveraging large-scale web text data during pre-training facilitates the acquisition of diverse linguistic patterns and nuances, thereby improving the models’ proficiency in both natural language comprehension and generation. Finally, the inclusion of media data offers the unique advantage of establishing an expansive and reliable media knowledge base. As shown in Fig. 3, we present a news-based training corpus where Chinese and Tibetan serve as the source languages, and English is provided as the translated language. Some of the original Chinese versions and Tibetan texts retain classical Chinese sentence structures, which are different from modern Chinese grammar.
@@ 3.1.2 Alpaca-Ti
Alpaca-Ti is used for instruction fine-tuning. We draw inspiration from Alpaca [31, 32] project at Stanford University and develop the first Tibetan benchmark specifically designed for self-supervised instruction fine-tuning in the world. It covers the following topics as shown in Table II.
TABLE II: Details of Alpaca-Ti
@@
Alpaca-Ti encompasses a diverse array of topics and also has a longer length and better instruction fine-tuning performance, designed to enhance LLM’s generalization capabilities. It has a total of 42,676 instances. One data sample is shown in Fig. 4.
Figure 4: One sample from Alpaca-Ti and ’-’ means nothing was entered.
@@ 3.1.3 Safety-Prompts-Ti
Safety-Prompts-Ti aims to improve the safety of large models on two major categories: Instruction Attacks and Typical Safety Scenarios, as shown in Table III, and the data size is 38,156. LLM’s representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts [22, 23, 24].
@@
TABLE III: Statistical Details of Instruction Attacks and Typical Safety Scenarios
Human evaluation is a valuable approach; however, due to the extensive workload it entails, we opted to create a benchmark to simulate the process in advance. One data sample is shown in Fig. 5.
Figure 5: One sample from Safety-Prompts-Ti.
@@ 3.1.4 CValues-Ti
CValues-Ti is the benchmark, used to improve the security of Tibetan language models. Considering work [25, 27, 28], we propose this branch of a small benchmark because LLM can pose risks or has negative social impacts. For CValues-Ti, it has three categories (safe & reponsibility = rejection & positive advice; safe = rejection; unsafe = risk response). Then, under the same prompt, various types of responses can be categorized into positive and negative samples, differentiated by their levels of difficulty, so that there are three common combinations (safe + unsafe; safe & reponsibility + unsafe; safe& reponsibility + safe). Unlike the work [28], We streamlined and condensed the instructions, and compared to their benchmark (20K), our sub-benchmark, CValues-Ti (2K), achieved equivalent results in subsequent experiments. One data sample is shown in Fig. 6.
Figure 6: One sample from CValues-Ti.
@@
hh-rlhf-Ti is also the first benchmark that improves the model effect by introducing human intervention. It consists of two parts: ”harmless (23,822)” and ”helpful (23,589)” just like RLHF [29, 30], but we reset and summarize the previous benchmark based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted prior to training for other objectives. This approach could be important in the future to prevent the development of intermediate, unaligned AI systems for Tibetan. It has a total of 47,411 instances, and one data sample is shown in Fig. 8.
@@ 3.2 Sun-Shine
Sun-Shine is built on the LLaMA-3 [2, 20] architecture, a collection of multilingual base models with 7 to 65 billion parameters, openly available to the research community. For more accessible training, we adopt the LLaMA 3.1 (8B) model [20]. Additionally, we integrate NLP model translation with human annotation [21] as supplementary data to enhance Sun-Shine’s accuracy.